# Systematic biases from the exclusion of higher harmonics in parameter estimation on LISA binaries

### v1

@@ VI.1.2 Approximate phases
$\Psi_{\ell m}\approx\frac{m}{2}\Psi_{22}\left(\frac{2f}{m}\right)+\Delta_{\ell m}+\left(1-\frac{m}{2}\right)\frac{\pi}{4}\,,$ (11)
with $\Delta_{\ell m}$ the inspiral-portion phase shift given in Eq. (7) of [52],
$\Delta_{\ell m}=\frac{\pi}{2}\left[3\ell+\mathrm{mod}(\ell+m,2)\right]-\pi\,.$ (12)
In Fig. 17, we plot the difference between the approximate phases computed via Eq. (11) and the “exact” phases from the IMRPhenomXHM waveform for the same $M=10^{6}\,M_{\odot},\,q=8$ system with amplitudes shown in Fig. 16. The approximate phases perform quite well at low frequencies and begin to veer off around the merger. We show the location of the merger frequency (i.e., the point at which the IMRPhenomXHM (3, 3) and (4, 4) amplitudes are at their maximum) with dotted lines.
As done for the amplitudes, we additionally observed whether higher-order phases might be better approximated from the (3, 3) or (4, 4) modes than the (2, 2), i.e., if we replace $m/2\times\Psi_{22}\left(2f/m\right)$ in Eq. (11) with $m/3\times\Psi_{33}\left(3f/m\right)$ or $m/4\times\Psi_{44}\left(4f/m\right)$. While we sometimes found that approximating from the (3, 3) resulted in a smaller difference from the “exact” phase and/or a more stable phase through merger, we only examined these phenomena for a handful of systems. For the present study, we build the phases simply following Eq. (11), and we leave a more thorough investigation of optimal ways to approximate higher-order modes for future study.
@@ VI.2 Biases with approximate waveforms
Figure 18: Parameter recovery with and without the extrapolated (5, 5) and (6, 6) modes (see Figs. 16 and 17). While virtually no bias is seen when excluding just the (6, 6) mode, there is a slight bias when excluding in addition the (5, 5) mode in the template for recovery. For reference, the extrapolated modes have SNR66=15.6 and SNR55=56, compared to SNR22=780.
@@
When injecting the signal with the extrapolated (5, 5) and (6, 6) modes and recovering without the (6, 6) (yellow contours and histograms), the bias is barely discernible. This is fairly unsurprising, given the low relative SNR of the (6, 6) mode (SNR66=15.6, compared to SNR55=56 and SNR22=780). When injecting with both extrapolated modes and recovering without either of them (purple contours and histograms), there are some biases. The recovered values of the intrinsic parameters are near enough to injected values that the bias is unlikely to make a significant difference. However, we note that a few extrinsic parameters, such as the ecliptic longitude and inclination, exhibit biases that could make some difference in the interpretation of the physical system.
In short, our results suggest that while the (5, 5) mode may increase slightly the accuracy of parameter inference of MBHBs, the (6, 6) mode is likely to make a much smaller impact. In addition to the difference in relative SNR of these higher-order modes, we note that in PN theory, even-$m$ modes are generally easier to reconstruct from other modes as compared to odd-$m$ modes. It is possible that when the $\ell\geq 5$ modes are removed, the remaining waveform has an easier time reproducing the (6, 6) than the (5, 5) mode, contributing to the lesser bias observed for this mode.
Finally, we note that the system we show these results for is one with some of the most “biased” results (high total mass and mass ratio, nearly edge-on). We aim to do a future study to determine the relative impact of (5, 5), (6, 6) and higher $\ell$ modes after marginalizing over inclination (for LISA and also for current and future ground based detectors) [54]. It is worth noting that we expect LISA to detect systems more massive than the ones we considered ($M>10^{6}\,M_{\odot}$), and we could also include such systems in our future study. Finally, the biases could be more pronounced for systems with spinning progenitor BHs, a consideration we do not address here.
@@ VII Conclusions
There are many avenues for future work. In the first place, it will eventually be important to understand how these biases would appear for binaries with spinning progenitors, including precessing systems. In fact, we expect higher-order modes to be more important for systems with high aligned spins (see, e.g., [26]). Similarly, we will eventually have to account for eccentric systems, systems with even higher mass ratios than what we have explored thus far ($q>10$), etc.
@@ Acknowledgements.
We thank Luca Reali, Veome Kapil, Rohit Chandramouli, Nicolás Yunes, Will Farr, and Leo Stein for helpful discussions. S. Y. is supported by the NSF Graduate Research Fellowship Program under Grant No. DGE2139757. S. Y., F. I., D. W, and E. B. are supported by NSF Grants No. AST-2307146, PHY-2207502, PHY-090003 and PHY-20043, by NASA Grants No. 20-LPS20-0011 and 21-ATP21-0010, by the John Templeton Foundation Grant 62840, by the Simons Foundation, and by the Italian Ministry of Foreign Affairs and International Cooperation grant No. PGR01167. The work of F. I. is supported by a Miller Postdoctoral Fellowship. S. M. acknowledges support form the French space agency CNES in the framework of LISA. This work was carried out at the Advanced Research Computing at Hopkins (ARCH) core facility (rockfish.jhu.edu), which is supported by the NSF Grant No. OAC-1920103.
@@
In this Appendix, we show how the use of different approximants (namely, IMRPhenomHM [52] and IMRPhenomXHM [42]) affects our PE results. Despite belonging to the same waveform family, these two approximants feature very different procedures to approximate the higher-order mode content of the signal, as we shall see below. Before we proceed, we note that to truly have a direct comparison between results with the different approximants, one would need to carefully align the waveforms in phase, which we do not do here. Our purpose is merely to show qualitatively the effect of using different methods to approximate higher-order modes.
In Fig. 19, we plot posteriors on intrinsic parameters as obtained in PE runs with IMRPhenomHM (left column) and IMRPhenomXHM (right column). For brevity, in this section we will drop the common “IMRPhenom” portion of each approximant name and simply use the labels “XHM” or “HM”. We show the comparisons for both a moderate-bias case ($M=3\times 10^{5}\,M_{\odot}$, top row) and a case with more severe biases ($M=10^{6}\,M_{\odot}$, bottom row).
In the moderate-bias case, the injected values of the intrinsic parameters are recovered fairly accurately with comparatively fewer modes in the template with HM. At first glance, this may be unexpected, since XHM is the newer approximant family (with all modes calibrated to numerical relativity simulations). However, we note that this is not necessarily indicative of HM performing better; on the contrary, it could be that the effect of higher-order modes is underestimated with this approximant, such that including more modes does not accurately capture the full extent of changes in the physical signal. This possibility is supported by the fact that XHM is much better than HM at recovering signals injected with NR surrogate waveforms (see, e.g., Fig. 14 of Ref. [42]).
@@
To summarize, we see that the systematic bias due to insufficient mode content does not appear in the same way when using different waveform approximants (with the former caveat that we are not looking at exactly the same systems, due to the phases not being exactly aligned). Importantly, in the HM family, the higher modes are more simply rescaled from the (2, 2) mode, indeed in much the same manner as we employ in Sec. VI of the main text. In contrast, each mode in the XHM family is calibrated to numerical relativity simulations, such that the modes are in a sense more truly independent from one another. Adding or removing one higher mode can therefore be more consequential with XHM than with HM (i.e., it can alter the waveform more significantly). This sort of phenomenon is important to keep in mind as we move towards more accurate waveform models containing higher-order modes.
@@ Appendix B Cutler-Vallisneri formalism
In this appendix, we outline for completeness the linear signal approximation introduced by Cutler and Vallisneri [19], which we use as a cheap method to approximate systematic biases due to insufficient mode content in waveform templates.
@@
$\mathrm{ln}\,p(s|\bm{\theta})\propto-\frac{1}{2}\left(s-h_{m}|s-h_{m}\right)\,$ (14)
@@
[…]
@@ Appendix C Additional results for the CV and direct likelihood optimization estimates of systematic biases
In Fig. 20, we plot biases as determined by the CV and direct likelihood optimization methods, as well as PE, for the remainder of the extrinsic parameters of the $M=10^{6}M_{\odot},\,q=8,\,\iota=\pi/12$ event discussed in the main text (Figs. 11 and 12).
Figure 20: Systematic bias on extrinsic parameters estimated within the linear signal approximation (“CV bias”) and with the direct optimization method (“likelihood bias”) compared to biases recovered in full PE (“ptemcee”). Results are shown for the $M=10^{6}M_{\odot},\,q=8,\,\iota=\pi/12$ event also discussed in Sec. V.

### v2

@@ VI.1.2 Approximate phases
$\Psi_{\ell m,\;\rm{approx}}(f)\approx\frac{m}{2}\Psi_{22}\left(\frac{2f}{m}\right)$ - $\Delta_{\ell m}$ (11)
- $\left(1-\frac{m}{2}\right)\frac{\pi}{4}\,$,
with $\Delta_{\ell m}$ given by
$\Delta_{\ell m}=\mathrm{mod}\left(\frac{m\pi}{2}-\pi,2\pi\right)$ (12)
[see, e.g., Ref. [Cotesta:2020qhw], although note the sign error in their Eq. (3.5)]. In Fig. 17, we plot the difference between the approximate phases computed via Eq. (11) and the “exact” phases from the IMRPhenomXHM waveform for the same $M=10^{6}\,M_{\odot},\,q=8$ system with amplitudes shown in Fig. 16. The approximate phases perform quite well at low frequencies and begin to veer off around the merger. We show the location of the merger frequency (i.e., the point at which the IMRPhenomXHM (3, 3) and (4, 4) amplitudes are at their maximum) with dotted lines.
As done for the amplitudes, we additionally observed whether higher-order mode phases might be better approximated from the (3, 3) or (4, 4) modes than the (2, 2), i.e., if we replace $m/2\times\Psi_{22}\left(2f/m\right)$ in Eq. (11) with $m/3\times\Psi_{33}\left(3f/m\right)$ or $m/4\times\Psi_{44}\left(4f/m\right)$. While we sometimes found that approximating from the (3, 3) resulted in a smaller difference from the “exact” phase and/or a more stable phase through merger, we only examined these phenomena for a handful of systems. For the present study, we build the phases simply following Eq. (11), and we leave a more thorough investigation of optimal ways to approximate higher-order modes for future study.
@@ VI.2 Biases with approximate waveforms
Figure 18: Parameter recovery with and without the extrapolated (5, 5) and (6, 6) modes (see Figs. 16 and 17). For reference, the extrapolated modes have SNR55=76.5 and SNR66=34.4, compared to SNR22=780.1. There are moderate biases when excluding the extrapolated modes from the template for recovery. While the biases are noticeable at this inclination angle near $\pi/2$, they are indiscernible for nearly face-on binaries, where the spin-weighted spherical harmonics for $(\ell,|m|)=(5,5)$ and $(6,6)$ are zero or very close to zero.
@@
When injecting the signal with the extrapolated (5, 5) and (6, 6) modes and recovering without the (6, 6) (yellow contours and histograms), we observe very slight biases in the recovery of both the intrinsic and extrinsic parameters. When removing the (5, 5) mode as well (purple contours), the biases on all parameters generally become somewhat worse.
We note that the relative importance of the (5, 5) and (6, 6) modes is heavily dependent upon the value of the inclination angle. In particular, at this inclination near $\pi/2$, the spin-weighted spherical harmonics $_{-2}Y_{55}(\iota,\phi=0.2)$ and $_{-2}Y_{66}(\iota,\phi=0.2)$ are larger than the harmonics for any other modes in our templates [i.e., for $(\ell,|m|)=$(2, 2), (3, 3), (2, 1), (4, 4), or (3, 2)]. The single-mode SNRs of the extrapolated modes at $\iota=\pi/2$ are SNR55=76.5 and SNR66=34.4, which are fairly comparable to the SNR of the (3, 2) mode at this inclination (SNR32=60.8); the total SNR of this system is 912.9. In contrast, at $\iota=\pi/12$ with all other parameters the same, we have SNR55=3.4 and SNR66=0.4 (compared to the total SNR=1769.4). The PE in this case would be virtually indistinguishable whether we exclude just the (6, 6) mode, both the (5, 5) and (6, 6), or neither extrapolated mode in the template for recovery.
In short, we find that depending on the inclination angle, waveforms containing only modes with $\ell\leq 4$ may not be sufficiently accurate to perform PE on MBHBs with total mass up to $10^{6}M_{\odot}$. Within the scope of this paper, we have only checked to see whether the (5, 5) and (6, 6) modes might be necessary within the parameter space under consideration. Eventually, we will need to determine how many modes are “enough”; will we also need $\ell\geq 7$ modes? How about $\ell>4$ modes with $\ell\neq|m|$? We plan to explore these questions in future work. Note also that the observable volume is different for systems with different inclinations (e.g., face-on systems have a larger observable volume compared to edge-on systems). In an upcoming study, we plan to marginalize over inclination including the observable volume effect. This will help us more accurately characterize the relative impact of higher-order modes for a system with given $M,q$ (for LISA and also for current and future ground-based detectors).Finally, the biases could be more pronounced for systems with spinning progenitor BHs, a consideration we do not address here.
@@ VII Conclusions
There are many avenues for future work. In the first place, it will eventually be important to understand how these biases would appear for binaries with spinning progenitors, including precessing systems. In fact, we expect higher-order modes to be more important for systems with high aligned spins (see, e.g., [Varma:2016dnf]). Similarly, we will eventually have to account for eccentric systems, systems with even higher mass ratios than what we have explored thus far ($q>10$), etc.
@@ Acknowledgements.
We thank Luca Reali, Veome Kapil, Rohit Chandramouli, Nicolás Yunes, Will Farr, and Leo Stein for helpful discussions. S.Y. is supported by the NSF Graduate Research Fellowship Program under Grant No. DGE2139757. S. Y., F. I., D. W, and E. B. are supported by NSF Grants No. AST-2307146, No. PHY-2513337, No. PHY-090003, and No. PHY-20043, by NASA Grant No. 21-ATP21-0010, by John Templeton Foundation Grant No. 62840, by the Simons Foundation [MPS-SIP-00001698, E.B.], by the Simons Foundation International [SFI-MPS-BH-00012593-02], and by Italian Ministry of Foreign Affairs and International Cooperation Grant No. PGR01167. The work of F.I. is supported by a Miller Postdoctoral Fellowship. S.M. acknowledges support form the French space agency CNES in the framework of LISA. This work was carried out at the Advanced Research Computing at Hopkins (ARCH) core facility (rockfish.jhu.edu), which is supported by the NSF Grant No. OAC-1920103.
@@
In this Appendix, we show how the use of different approximants (namely, IMRPhenomHM [London:2017bcn] and IMRPhenomXHM [Garcia-Quiros:2020qpx]) affects our PE results. Despite belonging to the same waveform family, these two approximants feature very different procedures to approximate the higher-order mode content of the signal, as we shall see below. Before we proceed, we note that to truly have a direct comparison between results with the different approximants, one would need to carefully align the waveforms in phase, which we do not do here. Our purpose is merely to show qualitatively the effect of using different methods to approximate higher-order modes.
In Fig. 19, we plot posteriors on intrinsic parameters as obtained in PE runs with IMRPhenomHM (left column) and IMRPhenomXHM (right column). For brevity, in this section we will drop the common “IMRPhenom” portion of each approximant name and simply use the labels “XHM” or “HM.” We show the comparisons for both a moderate-bias case ($M=3\times 10^{5}\,M_{\odot}$, top row) and a case with more severe biases ($M=10^{6}\,M_{\odot}$, bottom row).
In the moderate-bias case, the injected values of the intrinsic parameters are recovered fairly accurately with comparatively fewer modes in the template with HM. At first glance, this may be unexpected, since XHM is the newer approximant family (with all modes calibrated to numerical relativity simulations). However, we note that this is not necessarily indicative of HM performing better; on the contrary, it could be that the effect of higher-order modes is underestimated with this approximant, such that including more modes does not accurately capture the full extent of changes in the physical signal. This possibility is supported by the fact that XHM is much better than HM at recovering signals injected with NR surrogate waveforms (see, e.g., Fig. 14 of Ref. [Garcia-Quiros:2020qpx]). It is therefore likely that compared to XHM models, the older HM models do not adequately reflect the full dependence of the total waveform on higher-order mode content. A potential consequence of this would be that the HM models are therefore not as useful in understanding the full extent of biases due to neglecting higher-order modes.
@@
To summarize, we see that the systematic bias due to insufficient mode content does not appear in the same way when using different waveform approximants (with the former caveat that we are not looking at exactly the same systems, due to the phases not being exactly aligned). Importantly, in the HM family, the higher modes are more simply rescaled from the (2, 2) mode, indeed in much the same manner as we employ in Sec. VI of the main text. In contrast, each mode in the XHM family is calibrated to numerical relativity simulations, such that the modes are in a sense more truly independent from one another. Adding or removing one higher mode can therefore be more consequential with XHM than with HM (i.e., it can alter the waveform more significantly). This sort of phenomenon is important to keep in mind as we move toward more accurate waveform models containing higher-order modes.
@@ Appendix B Cutler-Vallisneri formalism
In this Appendix, we outline for completeness the linear signal approximation introduced by Cutler and Vallisneri [Cutler:2007mi], which we use as a cheap method to approximate systematic biases due to insufficient mode content in waveform templates.
@@
$\mathrm{ln}\,p(s|\bm{\theta})\propto-\frac{1}{2}\left(s-h_{m}|s-h_{m}\right)\,$, (14)
@@
$\left(a|b\right)=4\mathrm{Re}\int_{0}^{\infty}\frac{\tilde{a}(f)\tilde{b}^{*}(f)}{S_{n}(f)}df\,$, (16)
@@
Figure 20: Systematic bias on extrinsic parameters estimated within the linear signal approximation (“CV bias”) and with the direct likelihood optimization method (“likelihood bias”) compared to biases recovered in full PE (“ptemcee”). Results are shown for the $M=10^{6}M_{\odot},\,q=8,\,\iota=\pi/12$ event also discussed in Sec. V.
@@ Appendix C Additional results for the CV and direct likelihood optimization estimates of systematic biases
In Fig. 20, we plot biases as determined by the CV and direct likelihood optimization methods, as well as PE, for the remainder of the extrinsic parameters of the $M=10^{6}M_{\odot},\,q=8,\,\iota=\pi/12$ event discussed in the main text (Figs. 11 and 12). Both the CV and direct likelihood optimization methods recover the biases in PE fairly well for the more moderate biases observed with at least two modes in the waveform template, compared to the quadrupole-only case shown in Fig. 12.
Appendix D Effect of higher harmonics on the precision of sky localization estimates
In this Appendix, we expand on the fifth point of Sec. IV regarding the improved precision of the inference of sky position angles when using more than the (2, 2) mode. While the focus of the main text is on how parameter estimation on MBHBs becomes biased when higher harmonics are not adequately modeled in waveform templates, it is also true that the precision (in addition to the accuracy) of parameter estimation is increased considerably with the addition of higher-order modes.
To illustrate this point, in Fig. 21 we plot the maximum redshift at which an MBHB with the given total mass and mass ratio can be localized within $28\;\rm{deg}^{2}$ (i.e., the area in which GW170817 was localized) when using a quadrupole-only template (left panel), and then when adding one or more modes to the template (center and right panels, respectively). In these panels, the sky area $\Delta\Omega_{\rm{X}\%}$ at a confidence level $\rm{X}$ is computed as in Eq. (28) of Ref. [Iacovelli:2022bbs]:
$\Delta\Omega_{\rm{X}\%}$ $=$ $-2\pi|\cos\beta_{L}|\ln\left(1-\frac{\rm{X}}{100}\right)$
$\times$ $\sqrt{(\Gamma^{-1})_{\beta_{L}\beta_{L}}(\Gamma^{-1})_{\lambda_{L}\lambda_{L}}-(\Gamma^{-1})^{2}_{\beta_{L}\lambda_{L}}}\,$,
where $\Gamma$ is the Fisher information matrix and $(\Gamma^{-1})_{\theta^{i}\theta^{j}}$ denotes the component of the inverse Fisher (i.e., covariance) matrix for parameters $\theta^{i}$ and $\theta^{j}$.
Figure 21: Redshift at which MBHBs with the given total mass and mass ratio can be localized within 28 deg2 when using the (2, 2) mode only (left), the (2, 2) and (3, 3) modes (center), and the (2, 2), (3, 3), (2, 1) and (4, 4) modes (right). It is clear that sky localization improves considerably with just one additional mode beyond the (2, 2), while adding further modes beyond that does not introduce much further improvement at the masses we consider.
We stress that because we use Fisher matrices to compute the errors here, we are not accounting for the loss of precision due to the multimodal nature of the posteriors on $\lambda_{L}$ and $\beta_{L}$. While the posteriors on these parameters are for the most part unimodal in the parameter space considered within the scope of this paper (see Fig. 10), degeneracies in the sky position start to become more severe for more massive events, which last for a shorter amount of time. Moreover, in the (2, 2)-only case in particular, one can find apparent bias in the sky position posteriors that comes from projecting the multimodal posterior into the subspace of extrinsic parameters (see Eq. (79) of Ref. [Marsat:2020rtl]). We therefore caution the reader against taking the plots in Fig. 21 as perfectly confident estimates of the precision with which we will be able to localize MBHB events. Nevertheless, it is interesting to notice how dramatically the sky localization precision improves with the addition of even one mode beyond the (2, 2), whereas adding further modes beyond that does not make much of a difference at the masses we consider here. We point out that the (3, 2) mode is not included in the estimate of statistical error in the right panel of Fig. 21 because, as mentioned several times in the main text, this mode cannot be easily incorporated in the calculation of Fisher matrices due to differentiability issues.
Finally, we note that the sky localization depends heavily on the inclination angle. Results are shown in Fig. 21 for MBHB systems with a median inclination angle of $\iota=\pi/3$, but localization can improve significantly for more face-on binaries (and likewise worsen for more edge-on systems). The same plot made for $\iota=\pi/12$ results in an ability to localize systems within 28 deg2 out to redshift $z\sim 15$ when using four modes in the waveform.